Missions   Polar   Operations   Major Events

Polar MFE Major Events

The following is a list of major events that significantly affect the data the MFE produces.

The data saturates near perigee until the gain control by stored commands became active on 3/16/96.

The 9.2 sec sun crossing data is not valid until SPRP activation 3/15/96.

2002/5/22: An error in updating on-board offsets caused invalid returns of MFE data starting from 0050 UT of this date. The corrected on-board offsets were submitted to the spacecraft in an RTC interval on 2002/6/5.

Polar MFE Abnormal Data (2002/01/20-2002/02/01): Abnormal Polar MFE data flow occurred between 0100 UT on January 20, 2002 (DOY 020) and 1613 UT on February 1, 2002 (DOY 032). The incident was caused by an incorrect value stored during an activity load. This error affects only the storage of MFE data. The magnetic field vectors transferred to other instruments on board are not affected .

A series of RTC's were attempted to resolve the data flow problem. These RTC's resulted in several short data gaps and different types of abnormal data flow.

  • 2002/01/30: MFE instrument was shut down at 17:12:55 UT and restarted at 17:13:42 UT. The data flow appeared normal after the restart. However, due to the problem with transferring NRT data to UCLA we were not able to monitor the instrument response to subsequent commands. The data values obtained by POCC indicated that the data flow still had errors.
  • 2002/01/31: MFE instrument was shut down again between 21:01:55 and 21:04:36. The data flow after several command loads appeared to be incorrect. A review of the activity loads indicated that one value loaded for updating offset values was wrong.
  • 2002/02/01: A RTC attempt in the next NRT interval to correct the error. However, the MFE data flow did not became normal. It was realized later that POCC submitted an extra command that was not requested.
  • 2002/02/01: The set of values were stored to MFE again at 16:12:51 UT. The data flow became normal since then.

After this incident MFE commanding procedure has been modified to avoid this problem from recurring.


02/28/96 (DOY 060) Initial MFE activation Boom stowed no spin phase reference sensors rotated from their expected postions by ~ 30 degrees
03/05/96 (DOY 065) Mag Boom Deploy at 10:40 GMT.

Found x,z axes swapped from what was expected.


03/06/96 (DOY 066) 02:00 to 05:00 UT Flipper Checkout and Sensor Characterization test.

Outboard flipped both directions.

Inboard test aborted due to high temperature.


03/07/96 (DOY 067) 19:50 UT Dummy flip command sent to reorder x/z axes to proper positions. Suspect X axis is inverted.
03/14/96 (DOY 074) Despun platform activation
03/15/96 (DOY 075) SPRP activation Sun crossing data now valid.
03/16/96 (DOY 076) Stored Commands activated. Gain Changes to prevent saturation now active.
	At time:	    MFE gain changes to:                 
	perigee-02:20      Outboard/High Range (A2 State = Flip Right)
	perigee-01:00      Inboard /High Range (A2 state)
	perigee+01:00      Outboard/High Range (A2 state)
	perigee+03:00      Outboard/Low  Range (A2 state)

03/18/96 (DOY 78) Verified that only the X axis is inverted. Y and Z are properly phased. Started working on patch to correct problem
04/01/96 (DOY 092) MFEFLT1 patch successfully loaded and tested 17:35 UT.

Data properly phased until range change at 20:46 UT.


04/02/96 (DOY 093) 17:30 UT patch activated in range change activities.

DATA SHOULD BE PROPERLY PHASED EXCEPT DURING SENSOR CHARACTERIZATION TESTS!


04/11/96 (DOY 102) SENSOR CHARACTERIZATION TESTS

14:30 - 15:40 Inboard Sensor test
20:30 - 21:15 Outboard Sensor Test


05/28/96 (DOY 149) Started using the Inboard/Low Range from perigee+01:00 to perigee+03:00 and the Outboard/Hig Range from perigee-02:20 to perigee-01:00
	At time:	    MFE gain changes to:                 
	perigee-02:20      Outboard/High Range (A2 State)
	perigee-01:00      Inboard /High Range (A2 State)
	perigee+01:00      Inboard /Low  Range (A2 State)
	perigee+03:00      Outboard/Low  Range (A2 State)

07/02/96 (DOY 184) Started using the Inboard/Low Range from perigee-02:20 to perigee-01:00 and the Outboard/Hig Range from perigee+01:00 to perigee+03:00
	At time:	    MFE gain changes to:                 
	perigee-02:20      Inboard /Low  Range (A2 State)
	perigee-01:00      Inboard /High Range (A2 State)
	perigee+01:00      Outboard/high Range (A2 State)
	perigee+03:00      Outboard/Low  Range (A2 State)

08/28/96 (DOY 241) Onboard offset adjustment activated

By 22:02 UT, initial onboard offset adjustment was activated. Prior to 2202 UT the onboard offsets were set to zero. After 22:02 UT, following offsets (in counts) are being added onboard to correct the data.

                            X        Y       Z
 Inboard/High Range       -22       23      19
 Inboard/Low  Range      -179      178     113
Outboard/High Range        60      -49     -27
Outboard/Low  Range       500      -41    -105

09/19/96 (DOY 263) Outboard flipper activity 16:00-17:00 UT
10/02/96 (DOY 276) Starting from 17:32 UT, the sun crossing data are taking from 10 ms data stream. It reduces the jitter due timing and the phase lag due to onboard filtering.
10/08/96 (DOY 282) Outboard flipper activity 16:00-17:00 UT
10/11/96 (DOY 285) Inboard flipper activity 14:20-15:20 UT
10/17/96 (DOY 291) Updated onboard offset adjustment.

After 21:28 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range       -22       25     -17
         Inboard/Low  Range      -173      184    -117
        Outboard/High Range        46      -19      28
        Outboard/Low  Range       380     -165     270

10/23/96 (DOY 297) Outboard flipper activity 07:00-08:00 UT
10/24/96 (DOY 298) Inboard flipper activity 18:00-19:00 UT
11/08/96 (DOY 313) Flip the inboard sensor to the left position (A1 State) for its normal operation at 18:00 UT. The reason to do so is that the inboard sensor did not return to its proper position at A2 state after the 10/11/96 flipper test. We intend to leave the inboard sensor at its flip left (A1) state for the next 6 months.

After 18:00 UT when the inboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z became:

                                   X        Y       Z
         Inboard/High Range       -17       25     -22
         Inboard/Low  Range      -117      184    -173
        Outboard/High Range        46      -19      28
        Outboard/Low  Range       380     -165     270

Starting from DOY 313, the inboard sensor will be used only for 45K range (Perigee-01:00 to Perigee+01:00). The outboard sensor/high range will be used in 5K range. The new gain changes are:

	At time:	    MFE gain changes to:                 
	perigee-02:20      Outboard/High Range (A2 State)
	perigee-01:00      Inboard /High Range (A1 State = Flip Left)
	perigee+01:00      Outboard/high Range (A2 State)
	perigee+03:00      Outboard/Low  Range (A2 State)

11/08/96 (DOY 313) to 11/13/96 (DOY 318) The Z axis of the inboard sensor was inverted from normal after the inboard sensor was flipped to left position. The patch for the Z-axis inversion (to its normal direction) was activated on 11/13/96 DOY 318, 07:30 GMT.
11/20/96 (DOY 325) Updated onboard offset adjustment.

After 19:37 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range       -16       23     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        49      -16      28
        Outboard/Low  Range       394     -133     270


11/22/96 (DOY 327) Adjusted gain change times to make more use of the outboard sensor. The new gain change times are:
	At time:	    MFE gain changes to:                 
	perigee-02:20      Outboard/High Range (A2 State)
	perigee-00:50      Inboard /High Range (A1 State = Flip Left)
	perigee+00:40      Outboard/high Range (A2 State)
	perigee+03:00      Outboard/Low  Range (A2 State)

12/05/96 (DOY 340) Updated onboard offset adjustment (both inboard and outboard).

After 06:32 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range       -13       21      22
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        57       -9      28
        Outboard/Low  Range       458      -73     270

12/28/96 (DOY 363) Updated onboard offset adjustment (outboard only).

After 00:30 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range       -13       21     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        62       -5      28
        Outboard/Low  Range       495      -39     270

01/29/97 (DOY 029) Updated onboard offset adjustment (inboard only).

After 08:56 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range       -18       23     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        62       -5      28
        Outboard/Low  Range       495      -39     270

02/25/97 (DOY 056) Flip the outboard sensor to the left position (A1 State) for its normal operation at 18:00 UT.

After outboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range       -18       23     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        28       -5     -62
        Outboard/Low  Range       270      -39    -495

04/2/97 (DOY 092) Updated onboard offset adjustment (inboard and outboard).

After 10:05 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range       -14       21     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        31      -16     -62
        Outboard/Low  Range       281     -133    -495

04/22/97 (DOY 112) Flip the outboard sensor to the right position (A2 State) for its normal operation at 13:15 UT.

After outboard sensor started being operated at flip right (A2) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range       -14       21     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        31      -16      62
        Outboard/Low  Range       281     -133     495

05/27/97 (DOY 147) 4 hours 27 minutes burst mode test

The burst mode test was performed during 19:33:30-23:59:42. The 240 msec MFE data were stored in major frames 7654-9390.


05/28/97 (DOY 148) 1 hour 2 minutes burst mode test

The burst mode test was performed during 00:00:00-01:01:48. The 240 msec MFE data were stored in major frames 1- 405.


06/08/97 (DOY 159) Updated onboard offset adjustment (Outboard only)

After 09:00 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range       -14       21     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        58      -10      31
        Outboard/Low  Range       461      -83     281

06/10/97 (DOY 161) Flip the outboard sensor to the left position (A1 State) for its normal operation at 08:00 UT.

After outboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range       -14       21     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range        31      -10     -58
        Outboard/Low  Range       281      -83    -461

06/30/97 (DOY 181) Updated onboard offset adjustment (Both Outboard and Inboard)

After 01:03 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -016      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       034     -006    -061
        Outboard/Low  Range       302     -051    -497

07/05/97 (DOY 186) Flip the outboard sensor to the right position (A2 State) for its normal operation at 06:00 UT.

After outboard sensor started being operated at flip right (A2) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -016      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       061     -006     034
        Outboard/Low  Range       497     -051     302

07/18/97 (DOY 199) Updated onboard offset adjustment (Both Outboard and Inboard)

After 01:01 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -019      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       065     -006     036
        Outboard/Low  Range       519     -056     317


07/22/97 (DOY 203) Flip the outboard sensor to the left position (A1 State) for its normal operation at 06:00 UT.

After outboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -019      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       036     -006    -065
        Outboard/Low  Range       317     -056    -519

08/15/97 (DOY 227) Updated onboard offset adjustment (Both Outboard and Inboard)

After 17:00 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -016      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       034     -007    -065
        Outboard/Low  Range       304     -053    -519

08/20/97 (DOY 232) Flip the outboard sensor to the right position (A2 State) for its normal operation at 00:00 UT.

After outboard sensor started being operated at flip right (A2) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -016      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       065     -007     034
        Outboard/Low  Range       519     -053     304

09/26/97 (DOY 269) - 09/29/97 (DOY 272) At 269:19:30 UT on Sept 26 at the beginning of a real time pass it was found that the POLAR magnetometer (MFE) was not returning meaningful measurements. The playback data showed that the magnetometer stopped the normal operation at 970926 269:14:22 UT. At this time, the inboard sensor was physically in flip left position (A1) and the outboard sensor was physically in flip right postion (A2).

The instrument was turned off at 970926 269:23:44:45 UT and back on at 970926 269:23:56 UT and began taking good data in its default operating mode (inboard magnetometer/low range, with flip state undetermined).

However, when the magnetometer was commanded to switch to the outboard sensor by a later stored command, it reverted back to the default configuration again (inboard/low range).

After we reloaded all the flight software on 970929 272:19:47 UT, the operation became normal, and the magnetometer was able to take stored commands for magnetometer/gain changes.

For the whole period before the flight software reload (269:23:56 UT to 272:19:47 UT), the internal offset table was not activated, i.e., the the internal offsets corresponding to X, Y, and Z are:

                                   X        Y       Z
         Inboard/High Range        0        0       0 
         Inboard/Low  Range        0        0       0 
        Outboard/High Range        0        0       0 
        Outboard/Low  Range        0        0       0 

09/30/97 (DOY 273) Real time commanding in real time passes was used to properly set both inboard and outboard sensor axes and flip states.

By 273:00:12 UT on Sep 30, the correct onboard offset adjustment table was activated. The following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -014      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       051     -017     034
        Outboard/Low  Range       407     -141     304

10/14/97 (DOY 287) Updated onboard offset adjustment (Both Outboard and Inboard)

After 13:00 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -017      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       048     -023     034
        Outboard/Low  Range       378     -183     304

10/17/97 (DOY 290) Flip the outboard sensor to the left position (A1 State) for its normal operation at 05:10 UT.

After outboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -017      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       034     -023    -048 
        Outboard/Low  Range       304     -183    -378

11/04/97 (DOY 308) Updated onboard offset adjustment (Outboard Only)

After 04:00 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z

         Inboard/High Range      -017      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       037     -022    -048 
        Outboard/Low  Range       327     -176    -378

11/09/97 (DOY 313) Updated onboard offset adjustment (Both Outboard and Inboard)

After 01:43 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -014      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       032     -018    -048
        Outboard/Low  Range       287     -146    -378

11/22/97 (DOY 326) Updated onboard offset adjustment (Both Outboard and Inboard)

After 03:18 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       030     -011    -048
        Outboard/Low  Range       254     -094    -378

11/26/97 (DOY 330) 11 minutes burst mode test

The burst mode test was performed during 19:15:25-19:25:51. The 240 msec MFE data were stored in major frames 7537-7605.


12/02/97 (DOY 336) Flip the outboard sensor to the right position (A2 State) for its normal operation at 05:40 UT.

After outboard sensor started being operated at flip right (A2) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       048     -011     030
        Outboard/Low  Range       378     -094     254

12/08/97 (DOY 342) Updated onboard offset adjustment for outboard only.

After 01:01 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       062     -010     030
        Outboard/Low  Range       488     -074     254

12/23/97 (DOY 357) Flip the outboard sensor to the left position (A1 State) for its normal operation at 00:00 UT.

After outboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       030     -010    -062
        Outboard/Low  Range       254     -074    -488

01/06/98 (DOY 006) 21 minutes burst mode test

The burst mode test was performed during 18:08:57-18:29:40. The 240 msec MFE data were stored in major frames 7101-7236.


01/10/98 (DOY 010) Updated onboard offset adjustment for both inboard and outboard.

After 01:12 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -015      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       038     -007    -062
        Outboard/Low  Range       326     -051    -488

01/26/98 (DOY 026) 21 minutes burst mode test

The burst mode test was performed during 11:02:13-11:22:55. The 240 msec MFE data were stored in major frames 4319-4454.


1/29/98 (DOY 029) Flip the outboard sensor to the right position (A2 State) for its normal operation at 03:00 UT.

After outboard sensor started being operated at flip right (A2) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -015      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       062     -007     038
        Outboard/Low  Range       488     -051     326

02/15/98 (DOY 046) Updated onboard offset adjustment for outboard only.

After 01:01 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -015      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       066     -007     038
        Outboard/Low  Range       527     -051     326

02/21/98 (DOY 052) Flip the outboard sensor to the left position (A1 State) for its normal operation at 00:00 UT.

After outboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -015      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       038     -007    -066
        Outboard/Low  Range       326     -051    -527

03/07/98 (DOY 066) 42 minutes burst mode test

The burst mode test was performed during 09:52:15-10:33:31. The 240 msec MFE data were stored in major frames 3862-4131.


03/08/98 (DOY 067) Updated onboard offset adjustment for both inboard and outboard.

After 00:21 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       032     -010    -066
        Outboard/Low  Range       286     -076    -527

03/28/98 (DOY 087) Flip the outboard sensor to the right position (A2 State) for its normal operation at 08:00 UT.

After outboard sensor started being operated at flip right (A2) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       066     -010     032
        Outboard/Low  Range       527     -076     286

04/04/98 (DOY 094) Updated onboard offset adjustment for outboard sensors only.

After 01:01 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       048     -025     032
        Outboard/Low  Range       387     -201     286

04/19/98 (DOY 109) Flip the outboard sensor to the left position (A1 State) for its normal operation at 00:16 UT.

After outboard sensor started being operated at flip left (A1) state, the internal offsets corresponding to X, Y, and Z become:

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       032     -025    -048
        Outboard/Low  Range       286     -201    -387

05/01/98 (DOY 121) Bad data caused by bad telemetry for the period from 23:03 to 23:17 GMT.
05/06/98 (DOY 126) POLAR upset.

The Polar spacecraft and the magnetometer were affected by the X2 class flare occurred around 8 GMT, followed by a large proton event. Polar was affected at 10 GMT at which time the MFE telemetry received from the spacecraft became invalid. Time, Frame counts and data were all invalid.

At 13:22 the telemetry from the Mag returned to normal but the analog data was invalid. It appears that the digital processor was operating properly and the time commands were being recorded by the magnetometer.

At 15:57 the ADC power was commanded OFF and at 15:58 it was commanded ON. This procedure is necessary to reset the ADC protection circuitry and power up the ADC.

At this point all readings returned to nominal with the exception of the passive telemetry readings of the temperatures provided by the spacecraft.

The ADC temperature voltage started at 3.5V which is consistent with normal turn on sequencies at ~25Deg C.

After a warm up of 30 minutes this voltage was reduced to 3.18V and after a couple of hours the reading is 2.5V.


05/07/98 (DOY 127) Updated onboard offset adjustment for both the inboard and the outboard sensors.

After 01:31 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -010      020     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       034     -021    -048
        Outboard/Low  Range       303     -161    -387

05/18/98 (DOY 138) - 06/04/98 (DOY 155) On May 18, 1998, 00:00-00:16 GMT, the attempt to flip the outboard sensors from A1 postion to A2 postion was failed. Before 980518 00:00 GMT, the sensor was physically in A1 position, The flipper did not complete the flip, a 90 deg rotation of sensors 1 and 3. The data showed that the sensors stopped after only 5.7 deg roration. Thus the sensors were bascially still at A1 postion physically, but the onbard matrix used to map the data from Sensor coordinates to XYZ coordinates has changed to A2 state. This mapping caused the spin tone in MFE data showing in Y and Z axis.

During a NRT pass on 980602 (DOY 153), we corrected the onboard matrix that properly set the axes. The Z axis was still 5.7 deg from the spin axis.

During a NRT pass on 980604 (DOY 155), we flipped the sensors back to A1 postion by rotating 5.7 degree.


06/14/98 (DOY 165) Updated onboard offset adjustment for both the inboard and the outboard sensors.

After 03:09 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -014      020     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       033     -009    -048
        Outboard/Low  Range       299     -077    -387

07/20/98 (DOY 201) Updated onboard offset adjustment for both the inboard and the outboard sensors.

After 02:00 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -018      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       038     -008    -048
        Outboard/Low  Range       331     -072    -387

08/18/98 (DOY 230) Updated onboard offset adjustment for the inboard sensors only.

After 08:42 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -014      020     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       038     -008    -048
        Outboard/Low  Range       331     -072    -387

09/30/98 (DOY 273) POLAR anomaly. The ADC protection circuitry was activated and the ADC turned off. At 16:11 GMT, normal operation was resumed.

Following is a message from John Wainwright regarding the anomaly:

At approx. 14:00 UT yesterday (Sep 30, 1998, Day 273), an anomaly 
occurred in the POLAR GTM (GGS Telemetry Module). The anomaly was 
first observed at AOS (15:05 ). The symptoms were exactly those observed
in a similar event on May 6, 1998; that is: Loss of valid telemetry; MFE 
ADC upset; HYDRA eeprom write protect enabled (due to a loss of telemetry
sync.). PIXIE also showed an anomaly, though not an upset. The spacecraft
and affected instruments were restored to full operation by 16:00. There 
would be a loss of approx. 1 hr. 50 min. of data.  Replay of recorded 
data shows the anomaly occurring at 14:00:05. Proton and xray flux data 
plots from the SEC show an event occurring at approx. 14:00. 

At the time of the anomaly, the spacecraft was within umbra; exit from 
umbra was predicted at 14:12:27. In the May 6 event, the spacecraft 
was also in eclipse.

John Wainwright

10/03/98 (DOY 276) Updated onboard offset adjustment for the outboard sensors only.

After 01:10 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -014      020     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       030     -023    -048
        Outboard/Low  Range       276     -187    -387

10/11/98 (DOY 284) Updated onboard offset adjustment for both the inboard and the outboard sensors.

After 02:00 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -017      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       035     -028    -048
        Outboard/Low  Range       316     -217    -387

11/03/98 (DOY 307) POLAR MFE Interrupt Service Patch Activated

At 19:26 UT, the interrupt service patch was activated to reduce the frequency of noise spikes and the noise levels of the monitor voltages.


11/18/98 (DOY 322) Updated onboard offset adjustment for both the inboard and the outboard sensors.

After 02:00 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -019      025     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       024     -013    -048
        Outboard/Low  Range       218     -120    -387

11/27/98 (DOY 331) Updated onboard offset adjustment for both the inboard and the outboard sensors.

After 0200 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -011      018     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       026     -013    -048
        Outboard/Low  Range       238     -120    -387

12/22/98 (DOY 356) Updated onboard offset adjustment for both the inboard and the outboard sensors.

After 0140 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -013      020     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       029     -009    -048
        Outboard/Low  Range       257     -092    -387

01/14/99 (DOY 014) 11 minutes burst mode test

The burst mode test was performed during 18:32:59-18:43:25. The 240 msec MFE data were stored in major frames 7255-7323.


01/20/99 (DOY 020) 10 minutes burst mode test.

The bust mode was started at 0330 UT and it took about 4 hours to finish the output frames. During this interval the instrument was in 240 msec sampling mode. It would affect the onboard sun crossing time data.


01/20/99 (DOY 020) - 01/26/99 (DOY 026)

The instrument sampling rate was inadvertently left at 240 msec sampling mode because the instrument reset command was sent a couple of minutes too early before the end of burst mode output snapshot frames. It was corrected to 120 msec sampling mode during the RTC pass on 990126 15:15.

The onboard sun crossing time data were affected by the 240 msec sampling time.


02/08/99 (DOY 039) 4 hours 7 minutes burst mode test

The burst mode test was performed during 12:00:12-16:06:28. The 240 msec MFE data were stored in major frames 4695-6301.


03/01/99 (DOY 060) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:11:13-12:17:29. The 240 msec MFE data were stored in major frames 3203-4808.


03/09/99 (DOY 068) 4 hours 7 minutes burst mode test

The burst mode test was performed during 12:42:14-16:48:30. The 240 msec MFE data were stored in major frames 4973-6579.


03/23/99 (DOY 082) 4 hours 7 minutes burst mode test

The burst mode test was performed during 16:07:12-20:13:28. The 240 msec MFE data were stored in major frames 6309-7915.


03/24/99 (DOY 083) Updated onboard offset adjustment for outboard sensors only.

After 0648 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -013      020     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       028     -020    -048
        Outboard/Low  Range       244     -178    -387

04/03/99 (DOY 093) 4 hours 7 minutes burst mode test

The burst mode test was performed during 02:13:10-06:19:26. The 240 msec MFE data were stored in major frames 870-2476.


04/10/99 (DOY 100) 4 hours 7 minutes burst mode test

The burst mode test was performed during 13:20:18-17:26:34. The 240 msec MFE data were stored in major frames 5219-6825.


04/18/99 (DOY 108) 4 hours 7 minutes burst mode test

The burst mode test was performed during 18:37:11-22:43:26. The 240 msec MFE data were stored in major frames 7283-8889.


04/24/99 (DOY 114) 4 hours 7 minutes burst mode test

The burst mode test was performed during 18:06:14-22:12:29. The 240 msec MFE data were stored in major frames 7086-8692.


05/01/99 (DOY 121) 4 hours 7 minutes burst mode test

The burst mode test was performed during 12:57:14-17:03:30. The 240 msec MFE data were stored in major frames 5068-6674.


05/08/99 (DOY 128) 4 hours 7 minutes burst mode test

The burst mode test was performed during 07:01:11-11:07:26. The 240 msec MFE data were stored in major frames 2747-4353.


05/15/99 (DOY 135) 4 hours 7 minutes burst mode test

The burst mode test was performed during 17:54:12-22:00:28. The 240 msec MFE data were stored in major frames 7002-8606.


05/22/99 (DOY 142) 4 hours 7 minutes burst mode test

The burst mode test was performed during 11:13:13-15:19:29. The 240 msec MFE data were stored in major frames 4389-5995.


05/29/99 (DOY 149) 4 hours 7 minutes burst mode test

The burst mode test was performed during 05:21:18-09:27:34. The 240 msec MFE data were stored in major frames 2097-3703.


06/05/99 (DOY 156) 4 hours 7 minutes burst mode test

The burst mode test was performed during 16:07:16-20:13:32. The 240 msec MFE data were stored in major frames 6307-7913.


06/11/99 (DOY 162) Updated onboard offset adjustment for outboard sensors only.

After 0200 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -013      020     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       036     -016    -055
        Outboard/Low  Range       329     -154    -450

06/12/99 (DOY 163) 4 hours 7 minutes burst mode test

The burst mode test was performed during 09:30:16-13:36:32. The 240 msec MFE data were stored in major frames 3721-5327.


06/19/99 (DOY 170) 4 hours 7 minutes burst mode test

The burst mode test was performed during 03:40:11-07:46:27. The 240 msec MFE data were stored in major frames 1438-3042.


06/26/99 (DOY 177) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:20:11-18:26:26. The 240 msec MFE data were stored in major frames 5607-7213.


07/08/99 (DOY 189) Updated onboard offset adjustment for both inboard and outboard sensors.

After 0130 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -015      021     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       032     -008    -055
        Outboard/Low  Range       287     -099    -450

07/10/99 (DOY 191) 4 hours 7 minutes burst mode test

The burst mode test was performed during 02:00:18-06:06:34. The 240 msec MFE data were stored in major frames 786-2392.


07/19/99 (DOY 200) 4 hours 7 minutes burst mode test

The burst mode test was performed during 01:17:15-05:23:31. The 240 msec MFE data were stored in major frames 455-2060.


07/24/99 (DOY 205) 4 hours 7 minutes burst mode test

The burst mode test was performed during 06:04:13-10:10:29. The 240 msec MFE data were stored in major frames 2377-3983.


08/02/99 (DOY 214) - 08/10/99 (DOY 222) MFE Operations After August, 1999 Polar Spacecraft Anomaly

For details of this anomaly, please read the description by John Wainwright.

Overview

On August 3, 1999, Polar Operations staff found that the CAP was not responding to commands, etc, when Polar was interrogated. When reset, the spacecraft had turned off high voltages etc.

The spacecraft command table was lost in the upset. The re-start was scheduled for 990805 (DOY 217).

It appeared that MFE had a soft reset after the spacecraft upset. The internal offest table was reset to all zeros. The burst mode and 10 msec patches were deactivated and the memory used by these routines corrupted. The Instrument was left in the Inboard/Low Range state (default state).

The data quality was poor after the first reset which we now attribute to improper resychronization of the MFE software and spacecraft timing signals. The data quality went back to normal after the second soft reset on 980808.

990802 (DOY 214)

The spacecraft command system stopped shortly after MFE switching to the inboard/high range at 20:27:00 UT. MFE data were not available until after the CAP was reset.

990803 (DOY 215)

No MFE data were available before 07:36:03 UT.

At 07:36:03 UT, MFE data restarted with the instrument in Inboard/low range/flip left. MFE reached saturation from 15:13 to 15:52 UT around the perigee.

Note: The flip state showed as undetermined during this time. MFE treats this state as flip left which corresponded with the current physical state of the flipper.

990804 (DOY 216)

MFE continued working in its default state (inboard/low range/flip left) until 20:54 UT. MFE reached saturation from 09:02 - 09:44 UT around the perigee.

From 20:54 to 20:58 UT, RTC commanding was used to set the MFE to outboard/low range.

MFE operated at outboard/low range/flip left state since 20:58 UT.

990805 (DOY 217)

The stored command tables re-started at 00:00 UT.

The first gain change occurred at 00:16 UT when MFE switched from outboard/low range to outboard/high range.

At ~ 15:45 UT (DOY 217), the spacecraft started sending us the time information.

990806 (DOY 218)

At 19:58 UT, RTC commanding was used to restore MFE flip left state. Before the RTC, MFE was physically in flip left position with flip state undetermined to the spacecraft.

990807 (DOY 219)

From 00:59 to 01:04 UT, RTC commanding was used to restore the MFE internal offset table.

At 06:04 UT, the next gain change following the RTC, the offset table was activated.

990808 (DOY 220)

At 10:46 UT, an MFE soft reset was triggered by the burst mode test scheduled before the Polar upset since the burst mode patch had not been reloaded.

This soft reset caused the internal offset table resetting to all zeros and a default flip state.

After this soft reset, the resychronization of the MFE improved the quality of the data to normal. This indicates that the original reset/synchronization did not work properly. If the MFE is not properly synchronized with the spacecraft excess noise and data drops are expected.

990809 (DOY 221)

From 17:04 to 17:08 UT, RTC commanding was used to load and activate MFE inversion patch.

From 17:09 to 17:15 UT, RTC commanding was used to load and activate MFE 10 MSEC patch.

From 17:16 to 17:21 UT, RTC commanding was used to load and activate MFE interrupt patch.

From 18:16 to 18:19 UT, RTC commanding was used to restore the MFE internal offset table.

From 18:19 to 18:41 UT, RTC commanding was used to restore MFE flip left state.

Starting from 18:41 UT, the internal offset table and flip status were restored. MFE returned to normal operation.

990810 (DOY 222)

From 23:02 to 23:07 UT, RTC commanding was used to load the burst mode patch.


08/16/99 (DOY 228) 4 hours 7 minutes burst mode test

The burst mode test was performed during 10:03:16-14:09:31. The 240 msec MFE data were stored in major frames 3932-5538.


08/22/99 (DOY 234) 4 hours 7 minutes burst mode test

The burst mode test was performed during 09:33:14-13:39:30. The 240 msec MFE data were stored in major frames 3740-5346.


08/28/99 (DOY 240) 4 hours 7 minutes burst mode test

The burst mode test was performed during 09:03:12-13:09:28. The 240 msec MFE data were stored in major frames 3543-5149.


09/01/99 (DOY 244) Updated onboard offset adjustment for both inboard and outboard sensors.

After 0159 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -012      019     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       028     -013    -055
        Outboard/Low  Range       251     -121    -450

09/04/99 (DOY 247) 4 hours 7 minutes burst mode test

The burst mode test was performed during 02:40:10-06:46:25. The 240 msec MFE data were stored in major frames 1046-2652.


09/11/99 (DOY 254) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:50:10-18:56:25. The 240 msec MFE data were stored in major frames 5805-7411.


09/18/99 (DOY 261) 4 hours 7 minutes burst mode test

The burst mode test was performed during 07:17:11-11:23:27. The 240 msec MFE data were stored in major frames 2849-4455.


09/25/99 (DOY 268) 4 hours 7 minutes burst mode test

The burst mode test was performed during 00:59:12-05:05:28. The 240 msec MFE data were stored in major frames 388-1994.


10/02/99 (DOY 275) 4 hours 7 minutes burst mode test

The burst mode test was performed during 13:03:13-17:09:29. The 240 msec MFE data were stored in major frames 5107-6713.


10/09/99 (DOY 282) Updated onboard offset adjustment for both inboard and outboard sensors.

After 0304 UT, following offsets (in counts) are being added onboard to correct the data.

                                   X        Y       Z
         Inboard/High Range      -017      022     022
         Inboard/Low  Range      -107      182     173
        Outboard/High Range       036     -028    -055
        Outboard/Low  Range       327     -254    -450

10/12/99 (DOY 285) 3 hours 59 minutes burst mode test

The burst mode test was performed during 20:01:12-23:59:48. The 240 msec MFE data were stored in major frames 7227-8781.


10/23/99 (DOY 296) 4 hours 7 minutes burst mode test

The burst mode test was performed during 06:50:15-10:56:31. The 240 msec MFE data were stored in major frames 2677-4283.


11/05/99 (DOY 309) 4 hours 7 minutes burst mode test

The burst mode test was performed during 01:56:10-06:02:26. The 240 msec MFE data were stored in major frames 759-2364.


11/12/99 (DOY 316) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:45:16-18:51:31. The 240 msec MFE data were stored in major frames 5773-7378.


11/19/99 (DOY 323) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:40:09-12:46:25. The 240 msec MFE data were stored in major frames 3394-5000.


11/26/99 (DOY 330) 4 hours 7 minutes burst mode test

The burst mode test was performed during 02:00:14-06:06:30. The 240 msec MFE data were stored in major frames 786-2391.


12/06/99 (DOY 340) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:40:18-18:46:34. The 240 msec MFE data were stored in major frames 5743-7349.


12/12/99 (DOY 346) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:45:14-18:51:30. The 240 msec MFE data were stored in major frames 5761-7367.


12/25/99 (DOY 359) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:55:11-13:01:27. The 240 msec MFE data were stored in major frames 3492-5098.


01/12/00 (DOY 012) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:50:12-12:56:28. The 240 msec MFE data were stored in major frames 3458-5064.


01/19/00 (DOY 019) 4 hours 7 minutes burst mode test

The burst mode test was performed during 02:35:17-06:41:32. The 240 msec MFE data were stored in major frames 1014-2620.


01/26/00 (DOY 026) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:45:17-18:51:32. The 240 msec MFE data were stored in major frames 5775-7380.


02/02/00 (DOY 033) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:50:17-12:56:33. The 240 msec MFE data were stored in major frames 3459-5063.


02/14/00 (DOY 045) 4 hours 7 minutes burst mode test

The burst mode test was performed during 10:30:10-14:36:25. The 240 msec MFE data were stored in major frames 4111-5717.


02/21/00 (DOY 052) 4 hours 7 minutes burst mode test

The burst mode test was performed during 04:20:18-08:26:34. The 240 msec MFE data were stored in major frames 1699-3305.


03/09/00 (DOY 069) 4 hours 7 minutes burst mode test

The burst mode test was performed during 17:00:11-21:06:27. The 240 msec MFE data were stored in major frames 6655-8261.


03/16/00 (DOY 076) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:00:09-18:06:25. The 240 msec MFE data were stored in major frames 5479-7083.


03/23/00 (DOY 083) 4 hours 7 minutes burst mode test

The burst mode test was performed during 11:20:13-15:26:28. The 240 msec MFE data were stored in major frames 4438-6044.


04/09/00 (DOY 100) 4 hours 7 minutes burst mode test

The burst mode test was performed during 11:30:18-15:36:34. The 240 msec MFE data were stored in major frames 4502-6108.


04/23/00 (DOY 114) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:00:11-12:06:27. The 240 msec MFE data were stored in major frames 3133-4737.


04/30/00 (DOY 121) 4 hours 7 minutes burst mode test

The burst mode test was performed during 15:00:18-19:06:34. The 240 msec MFE data were stored in major frames 5873-7479.


05/07/00 (DOY 128) 4 hours 7 minutes burst mode test

The burst mode test was performed during 09:30:09-13:36:25. The 240 msec MFE data were stored in major frames 3718-5324.


05/17/00 (DOY 138) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:00:18-12:06:33. The 240 msec MFE data were stored in major frames 3133-4739.


05/24/00 (DOY 145) 4 hours 7 minutes burst mode test

The burst mode test was performed during 18:00:16-22:06:32. The 240 msec MFE data were stored in major frames 7044-8650.


06/01/00 (DOY 153) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:00:12-12:06:28. The 240 msec MFE data were stored in major frames 3133-4739.


06/24/00 (DOY 176) 4 hours 7 minutes burst mode test

The burst mode test was performed during 04:10:12-08:16:28. The 240 msec MFE data were stored in major frames 1633-3239.


07/01/00 (DOY 183) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:00:11-12:06:26. The 240 msec MFE data were stored in major frames 3133-4738.


07/08/00 (DOY 190) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:40:12-18:46:28. The 240 msec MFE data were stored in major frames 5741-7347.


08/01/00 (DOY 214) 4 hours 7 minutes burst mode test

The burst mode test was performed during 16:10:10-20:16:25. The 240 msec MFE data were stored in major frames 6329-7935.


08/09/00 (DOY 222) 4 hours 7 minutes burst mode test

The burst mode test was performed during 04:30:17-08:36:33. The 240 msec MFE data were stored in major frames 1761-3367.


08/16/00 (DOY 229) 4 hours 7 minutes burst mode test

The burst mode test was performed during 17:40:14-21:46:30. The 240 msec MFE data were stored in major frames 6914-8520.


08/23/00 (DOY 236) 4 hours 7 minutes burst mode test

The burst mode test was performed during 12:04:15-16:10:31. The 240 msec MFE data were stored in major frames 4721-6325.


08/30/00 (DOY 243) 4 hours 7 minutes burst mode test

The burst mode test was performed during 06:40:15-10:46:30. The 240 msec MFE data were stored in major frames 2612-4218.


09/16/00 (DOY 260) 4 hours 7 minutes burst mode test

The burst mode test was performed during 14:55:10-19:01:26. The 240 msec MFE data were stored in major frames 5840-7445.


09/22/2000 (DOY 266)- 09/26/0000 (DOY 270) MFE Operation after Septemtember, 2000 Polar Spacecraft Anomaly

The Polar MFE instrument is operating normally after reloading software patches and reinitializing the flipper control circuits. This activity occured between 16:00 and 17:00 on day 270, 2000

In the past, the MFE instrument has suffered a ADC reset when there was an associated Polar GTM anomaly from which we could easily recover by cycling the ADC power and restoring the onboard high rate data. This was attempted on 266 at 23:15. The nominal data looked reasonable and the highrate data was being received.

In this case, however, some of the MFE software patches were also affected and when the normally scheduled gain change commands were executed and the patches activated, the instrument reset itself.

This explains why the highrate data was reset to off over the weekend and when we tried to reset the flipper status last night. Thus the initial GTM anomaly explains all the anomylous operations seen over the weekend.

It appears that all patches were loaded properly and execute as designed by 17:00.

There is some confusion about whether the gain/range command is working properly and this is under investigation.

We will continue to monitor the instrument to be sure that the scheduled commands properly execute.


10/27/00 (DOY 301) 4 hours 7 minutes burst mode test

The burst mode test was performed during 11:10:13-15:16:29. The 240 msec MFE data were stored in major frames 4372-5978.


11/03/00 (DOY 308) 4 hours 7 minutes burst mode test

The burst mode test was performed during 13:00:17-17:06:33. The 240 msec MFE data were stored in major frames 5090-6696.


11/10/00 (DOY 315) 4 hours 7 minutes burst mode test

The burst mode test was performed during 07:10:12-11:16:28. The 240 msec MFE data were stored in major frames 2807-4413.


11/18/00 (DOY 323) 4 hours 7 minutes burst mode test

The burst mode test was performed during 16:31:11-20:37:27. The 240 msec MFE data were stored in major frames 6464-8070.


11/25/00 (DOY 330) 4 hours 7 minutes burst mode test

The burst mode test was performed during 11:49:11-15:55:27. The 240 msec MFE data were stored in major frames 4627-6231.


12/02/00 (DOY 337) 4 hours 7 minutes burst mode test

The burst mode test was performed during 07:28:11-11:34:27. The 240 msec MFE data were stored in major frames 2924-4530.


12/10/00 (DOY 345) 4 hours 7 minutes burst mode test

The burst mode test was performed during 15:55:12-20:01:27. The 240 msec MFE data were stored in major frames 6223-7828.


12/17/00 (DOY 352) 4 hours 7 minutes burst mode test

The burst mode test was performed during 12:28:11-16:34:26. The 240 msec MFE data were stored in major frames 4881-6487.


12/24/00 (DOY 359) 4 hours 7 minutes burst mode test

The burst mode test was performed during 07:46:10-11:52:26. The 240 msec MFE data were stored in major frames 3042-4646.


01/06/01 (DOY 006) 4 hours 7 minutes burst mode test

The burst mode test was performed during 17:00:10-21:06:26. The 240 msec MFE data were stored in major frames 6655-8260.


01/14/01 (DOY 014) 4 hours 7 minutes burst mode test

The burst mode test was performed during 08:50:12-12:56:28. The 240 msec MFE data were stored in major frames 3456-5062.


01/21/01 (DOY 021) 4 hours 7 minutes burst mode test

The burst mode test was performed during 04:32:17-08:38:32. The 240 msec MFE data were stored in major frames 1777-3383.


01/28/01 (DOY 028) 4 hours 7 minutes burst mode test

The burst mode test was performed during 17:50:12-21:56:28. The 240 msec MFE data were stored in major frames 6977-8583.


02/10/01 (DOY 041) 4 hours 7 minutes burst mode test

The burst mode test was performed during 13:50:14-17:56:30. The 240 msec MFE data were stored in major frames 5414-7020.


02/19/01 (DOY 050) 4 hours 7 minutes burst mode test

The burst mode test was performed during 16:00:18-20:06:34. The 240 msec MFE data were stored in major frames 6261-7867.


02/25/01 (DOY 056) 4 hours 7 minutes burst mode test

The burst mode test was performed during 06:25:11-10:31:26. The 240 msec MFE data were stored in major frames 2514-4119.


06/16/01 (DOY 167) Updated onboard offset adjustment for both inboard and outboard sensors.

After 0042 UT, following offsets (in counts) are being added onboard to correct the data.

                              X        Y       Z
  	 Inboard/High Range      -012      020     022
	 Inboard/Low  Range      -107      182     173
	Outboard/High Range       028     -008    -055
	Outboard/Low  Range       256     -086    -450

09/15/01 (DOY 258) - 10/03/01 (DOY 276) Flip maneuvers were executed on September 15-16 and on October 1-3. Polar spacecraft also experienced a series of eclipse periods between September 17 and October 2.





Visitor - Research - Library - Education - Missions - Data - Personnel - Resources - News - Help

For more information about Space Science Center contact: www@igpp.ucla.edu.

Last updated: September 25, 2001